# Privacy-Preserving Neural Graph Databases

### 19 Jan 2024 v2

@@ 3.2. Problem Formulation
Suppose that there is a graph $\mathcal{G}=(\mathcal{V},\mathcal{R},\mathcal{A},\mathcal{N})$, where part of entities’ attributes are regarded as private information $\mathcal{A}=\mathcal{A}_{\text{private}}\cup\mathcal{A}_{\text{public}}$, where $a_{p}\in\mathcal{A}_{\text{private}}$ denotes the entity $u$ has the sensitive attribute $a_{p}$ with value $x$ if $a_{p}(u,x)=1$. Due to the variation in privacy requirements, an attribute can be handled differently for various entities. NGDBs store the graph in embedding space and can be queried with arbitrary complex logical queries. Due to the generalization ability of NGDBs, attackers can easily infer sensitive attributes $\mathcal{A}_{\text{private}}$ utilizing complex queries. To protect private information, NGDBs should retrieve obfuscated answers when queried with privacy risks while preserving high accuracy in querying public information.
@@ 4.1. Computational Graph
  • Projection: Projection has various types. Given a set of entities $S$, a relation between entities $r\in\mathcal{R}$, an attribute type $a$, a set of values $S\subset\mathcal{N}$, projection describes the entities or values can be achieved under specified relation and attribute types. For example, attribute projection is denoted as $P_{a}(S)=\{x\in\mathcal{N}|v\in S,a(v,x)=1\}$ describing the values that can be achieved under the attribute type $a$ by any of the entities in $S$. Furthermore, we have relational projection $P_{r}(S)$, reverse attribute projection $P_{a}^{-1}(S)$ and numerical projection $P_{f}(S)$. We denote all these projections as $P(S)$;
@@
  • Union: Given a set of entity or value sets $S_{1},S_{2},...,S_{n}$, the operator computes the union of $\cup_{i=1}^{n}S_{i}$
@@ 4.2. Privacy Threatening Query Answers
After a complex query is parsed into the computational graph $G$, and $M=G(S)$ denotes the answer set of entities or values retrieved by the computational graph. We divide the $M$ retrieved by the complex queries as private answer set $M_{\text{private}}$ and non-private answer set $M_{\text{public}}$ based on whether the answer has to be inferred under the involvement of private information. As shown in Figure 2, query in (A) retrieves answers from the knowledge graph, Toronto is regarded as privacy-threatening answers because $LiveIn(Hinton,Toronto)\in\mathcal{A}$.
@@
For the projection on private inputs,
@@
As shown in Figure 3(A), all the red nodes within the answer sets can solely be deduced from private attribute links and are classified as query answers posing privacy risks. On the other hand, nodes such as the green-red node, while accessible through private projection, can also be inferred from public components, thereby carrying lower privacy risks.
Figure 2. An example of query demonstrating the retrieved privacy-threatening query answers. Red node denotes privacy risks. (A) The logic knowledge graph query involves privacy information. (B) Example of a complex query in the knowledge graph, Toronto is regarded as a privacy-threatening answer as it has to be inferred by sensitive information.
Figure 3. Example of privacy-threatening answer sets computation in projection, intersection, and union. Green nodes denote non-private answers, red nodes denote privacy-threatening answers, green-red nodes denote different privacy risks in subsets. Red dashed arrows denote privacy projection. The range demarcated by the red dashed lines denotes privacy-threatening answer sets.
@@
5. Privacy-Preserving Neural Graph Database
In this section, we present the privacy-preserving neural graph databases for protecting sensitive information in knowledge graphs while preserving high-quality complex query answering performance. We set two optimization goals for P-NGDBs: preserve the accuracy of retrieving non-private answers and obfuscate the privacy-threatening answers.
@@ 5.1. Encoding representation
The query encoding methods encode queries to embedding space and compare the similarity with the entities or attributes for answer retrieval, the encoding process can be represented as:
(2) $q=f_{g}(Query),$
where $f_{g}$ is the query encoding methods parameterized by $\theta_{g}$, and $q\in R^{d}$ represents the query embedding for simplicity. Given a query, we iteratively compute the queries based on the sub-query embeddings and logical operators. Assume the encoding process in step $i$, the subquery embedding is $q_{i}$. We can denote the logical operators projection, intersection, and union as:
@@
where $f_{p}$, $f_{i}$, and $f_{u}$ denote the parameterized projection, intersection, and union operators, respectively. After query encoding, we compute the score for every candidate based on the query encoding and entity (attribute) embeddings. Finally, we calculate the normalized probability using the Softmax function:
@@ 5.2. Learning Objective
In privacy-preserving neural graph databases, there are two learning objectives: given a query, P-NGDBs should accurately retrieve non-private answers and obfuscate the private answers. For public retrieval, given query embedding $q$, the loss function can be expressed as:
@@ 6. Experiments
In this section, to evaluate the privacy leakage problem in NGDBs and the protection ability of our proposed Privacy-preserving Neural Graph Database (P-NGDBs), we construct a benchmark on three real-world datasets and evaluate P-NGDB’s performance based on that.
@@ 6.1. Datasets
We create our benchmarks based on the numerical complex query datasets proposed in (Bai et al., 2023a), which is constructed on FB15K-numerical, DB15K-numerical, and YAGO15K-numerical (Kotnis and García-Durán, 2018), three publicly available knowledge graphs datasets. In each knowledge graph, there are vertices describing entities and attributes and edges describing entity relations, entity attributes, and numerical relations. We first randomly select a set of edges denoting entities’ attributes as privacy information. Then We divide the remaining edges with a ratio of 8:1:1 to construct training, validation, and testing edge sets, respectively. Then the training graph $\mathcal{G}_{train}$, validation graph $\mathcal{G}_{val}$, and testing graph $\mathcal{G}_{test}$ are constructed on training edges, training+validation edges and training+validation+testing edges, respectively. The detailed statistics of the three knowledge graphs are shown in Table 1, #Nodes represents the number of entities and attributes. #Edeges represents the number of relation triples and attribute triples. #Pri. Edges represents the number of attribute triples that are considered private.
@@ 6.2. Benchmark Construction
Following previous work (Hamilton et al., 2018; Bai et al., 2023a), we evaluate the complex query answering performance on the following eight general query types with abbreviations of $1p,2p,2i,3i,ip,ip,2u$, and $up$, as shown in Figure 4. For each general query type, each edge represents either a projection or a logical operator, and each node represents either a set of entities or numerical values. We use the sampling method proposed in (Bai et al., 2023a) to randomly sample complex queries from knowledge graphs. We random sample training, validation, and testing queries from the formerly constructed graphs respectively. For the training queries, we search for corresponding training answers on the training graph. For the validation queries, we use those queries that have different answers on the validation graph from answers on the training graph to evaluate the generalization ability. For the testing queries, we conduct a graph search on the testing graph with private edges for testing answers and we split the answers into private set and public set according to the definition of privacy risk query answers discussed in Section 4. We conduct a statistical analysis of the number of privacy risk answers for different types of complex queries of three knowledge graphs and the statistics are shown in Table 2.
@@ 6.3.1. Baselines
. Our proposed P-NGDBs can be applied to various complex query encoding methods to provide privacy protections. We select three commonly used complex query encoding methods and compare the performance with and without P-NGDB’s protection.
@@ 6.3.2. Evaluation Metrics
The evaluation consists of two distinct parts: reasoning performance evaluation and privacy protection evaluation. In performance evaluation, following the previous work (Bai et al., 2023a), we evaluate the generalization capability of models by calculating the rankings of answers that cannot be directly retrieved from an observed knowledge graph. Given a testing query $q$, the training, validation, and public testing answers are denoted as $M_{train}$, $M_{val}$, and $M_{test}$, respectively. We evaluate the quality of retrieved answers using Hit ratio (HR) and Mean reciprocal rank (MRR). HR@K metric evaluates the accuracy of retrieval by measuring the percentage of correct hits among the top K retrieved items. The MRR metric evaluates the performance of a ranking model by computing the average reciprocal rank of the first relevant item in a ranked list of results. The metric can be defined as:
(11) $\text{Metric}(q)=\frac{1}{|M_{test}/M_{val}|}\sum_{v\in M_{test}/M_{val}}m(rank(v)),\$
where $m(r)=\textbf{1}[r\leq K]$ if the metric is HR@K and $m(r)=\frac{1}{r}$ if the metric is MRR. Higher values denote better reasoning performance. In privacy protection evaluation, we compute the rankings metric of privacy-threatening answers as these answers cannot be inferred from the observed graphs, and smaller values denote stronger protection. We train all the models by using the training queries and private attributes and tune hyper-parameters using the validation queries. The evaluation is then finally conducted on the testing queries, including the evaluation of public answers for performance assessment and the evaluation of private answers for privacy assessment. The experiment results are reported on the testing public and private queries respectively.
@@
Table 4. The mean reciprocal ranking (MRR) variation for P-NGDBs on different types of queries.
@@ 6.5. Sensitivity Study
The privacy protection needs can be different in various knowledge graphs. We evaluate the impact of the privacy penalty coefficient $\beta$, A larger $\beta$ can provide stronger privacy protection. We select $\beta$ from $\{0.01,0.05,0.1,0.5,1\}$, and evaluate the retrieval accuracy of public answers and private answers, respectively. We evaluate GQE’s retrieval performance on the FB15K-N dataset and depict the MRR’s change under each $\beta$ compared to the unprotected GQE model. The results are shown in Figure 5. We can know that the penalty can effectively control the level of privacy protection, to meet higher privacy needs, P-NGDBs have to undertake much utility loss. Besides, the flexible adjustment of the privacy coefficient can make P-NGDBs adapt to various scenarios with different privacy requirements.

### 22 Feb 2024 v3

@@ 3.2. Problem Formulation
Suppose that there is a graph $\mathcal{G}=(\mathcal{V},\mathcal{R},\mathcal{A})$, where part of entities’ attributes are regarded as private information $\mathcal{A}=\mathcal{A}_{\text{private}}\cup\mathcal{A}_{\text{public}}$, where $a_{p}\in\mathcal{A}_{\text{private}}$ denotes the entity $u$ has the sensitive attribute $a_{p}$ with value $x$ if $a_{p}(u,x)=1$. Due to the variation in privacy requirements, an attribute can be handled differently for various entities. NGDBs store the graph in embedding space and can be queried with arbitrary complex logical queries. Due to the generalization ability of NGDBs, attackers can easily infer sensitive attributes $\mathcal{A}_{\text{private}}$ utilizing complex queries. To protect private information, NGDBs should retrieve obfuscated answers when queried with privacy risks while preserving high accuracy in querying public information.
@@ 4. Innocuous yet Private Query
Figure 2. An example of query demonstrating the retrieved privacy-threatening query answers. The red node denotes privacy risks. (A) The logic knowledge graph query involves privacy information. (B) An example of a complex query in the knowledge graph. Toronto is regarded as a privacy-threatening answer as it has to be inferred by sensitive information.
Figure 3. Example of privacy-threatening answer sets computation in projection, intersection, and union. Green nodes denote non-private answers, orange nodes denote privacy-threatening answers, and orange-green nodes denote different privacy risks in subsets. Red dashed arrows denote privacy projection. The range demarcated by the red dashed lines denotes privacy-threatening answer sets. The answers circled in red dashed line are at risk to leak privacy.
@@ 4.1. Computational Graph
  • Projection: Projection has various types for attribute- and value-sensitive queries (Bai et al., 2023a). Given a set of entities $S$, a relation between entities $r\in\mathcal{R}$, an attribute type $a$, a set of values $S\subset\mathcal{N}$, projection describes the entities or values can be achieved under specified relation and attribute types. For example, attribute projection is denoted as $P_{a}(S)=\{x\in\mathcal{N}|v\in S,a(v,x)=1\}$ describing the values that can be achieved under the attribute type $a$ by any of the entities in $S$. Furthermore, we have relational projection $P_{r}(S)$, reverse attribute projection $P_{a}^{-1}(S)$ and numerical projection $P_{f}(S)$. We denote all these projections as $P(S)$;
@@
  • Union: Given a set of entity or value sets $S_{1},S_{2},...,S_{n}$, the operator computes the union of $\cup_{i=1}^{n}S_{i}$.
@@ 4.2. Privacy Threatening Query Answers
After a complex query is parsed into the computational graph $G$, and $M=G(S)$ denotes the answer set of entities or values retrieved by the computational graph. We divide the $M$ retrieved by the complex queries as private answer set $M_{\text{private}}$ and non-private answer set $M_{\text{public}}$ based on whether the answer has to be inferred under the involvement of private information. As shown in Figure 2, query in (A) retrieves answers from the knowledge graph, Toronto is regarded as a privacy-threatening answer because the answer Toronto denotes the private information $LiveIn(Hinton,Toronto)$. To analyze the risks, we define private answer sets for different operators as follows. We denote the input set of operators as $S=S_{\text{private}}\cup S_{\text{public}}$, where $S_{\text{private}}$, $S_{\text{public}}$ denote whether the elements in the set involve private information.
@@
If an answer can only be inferred from private information, it will be regarded as a privacy threatening answer. For the projection on private inputs,
@@
The answers of projection on private inputs are privacy threatening answers as they can expose the existence of private inputs. As shown in Figure 3(A), all the red nodes within the answer sets can be solely deduced from private attribute links and are classified as query answers posing privacy risks. On the other hand, nodes such as the green-red node, while accessible through private projection, can also be inferred from public components, thereby carrying lower privacy risks.
@@
5. Privacy-Preserved Neural Graph Database
In this section, we mainly focus on the query encoding methods in NGDBs and present the privacy-preserved neural graph databases for protecting sensitive information in knowledge graphs while preserving high-quality complex query answering performance. We set two optimization goals for P-NGDBs: preserve the accuracy of retrieving non-private answers and obfuscate the privacy-threatening answers.
@@ 5.1. Encoding Representation
The query encoding methods can be decomposed into two steps: encode queries to embedding space $q\in\mathbb{R}^{d}$ and compare the similarity with the entities or attributes for answer retrieval. Given a query, we iteratively compute the queries based on the sub-query embeddings and logical operators. Assume the encoding process in step $i$, the sub-query embedding is $q_{i}$. We can denote the logical operators projection, intersection, and union as:
@@
where $f_{P}$, $f_{I}$, and $f_{U}$ denote the parameterized projection, intersection, and union operators, respectively. Here, the $f_{P}$, $f_{I}$, and $f_{U}$ can be instantiated with various types of parameterization of query encoding functions. We list all the instantiations used in our experiments in the Appendix A. After query encoding, we compute the score for every candidate based on the query encoding and entity (attribute) embeddings. Finally, we calculate the normalized probability using the Softmax function:
@@ 5.2. Learning Objective
In privacy-preserved NGDBs, there are two learning objectives: given a query, P-NGDBs should accurately retrieve non-private answers and obfuscate the private answers. For public retrieval, given query embedding $q$, the loss function can be expressed as:
@@ 6. Experiments
In this section, to evaluate the privacy leakage problem in NGDBs and the protection ability of our proposed Privacy-preserved Neural Graph Database (P-NGDBs), we construct a benchmark on three real-world datasets and evaluate P-NGDB’s performance based on that.
@@ 6.1. Datasets
There are various knowledge graph datasets used in complex query answering tasks (Bollacker et al., 2008; Carlson et al., 2010). Without loss of generality, we consider the privacy problem in the three commonly used multi-relational knowledge graph with numerical attributes: FB15K-N, DB15K-N, and YAGO15K-N (Kotnis and García-Durán, 2018) for several reasons: First, attribute value projections can be the same as traditional relation projection if the values themselves are entities, e.g., locations; Second, attributes and their values are more aligned with real-world privacy considerations; Third, attribute values are vulnerable to be attacked as we can use group queries to attack individual’s information, which has been widely used as an illustration in differential privacy (Dwork, 2008). We create our benchmarks based on the numerical complex query datasets proposed in (Bai et al., 2023a). In each knowledge graph, there are vertices describing entities and attributes and edges describing entity relations, entity attributes, and numerical relations. We first randomly select a set of edges denoting entities’ attributes as privacy information. Then We divide the remaining edges with a ratio of 8:1:1 to construct training, validation, and testing edge sets, respectively. Then the training graph $\mathcal{G}_{train}$, validation graph $\mathcal{G}_{val}$, and testing graph $\mathcal{G}_{test}$ are constructed on training edges, training+validation edges and training+validation+testing edges, respectively. The detailed statistics of the three knowledge graphs are shown in Table 1, #Nodes represents the number of entities and attributes. #Edeges represents the number of relation triples and attribute triples. #Pri. Edges represent the number of attribute triples that are considered private.
@@ 6.2. Benchmark Construction
Following previous work (Hamilton et al., 2018; Bai et al., 2023a; Wang et al., 2021b), we evaluate the complex query answering performance on the following eight general query types with abbreviations of $1p,2p,2i,3i,ip,pi,2u$, and $up$, as shown in Figure 4. For each general query type, each edge represents either a projection or a logical operator, and each node represents either a set of entities or numerical values. We use the sampling method proposed in (Bai et al., 2023a) to randomly sample complex queries from knowledge graphs. We random sample training, validation, and testing queries from the formerly constructed graphs respectively. For the training queries, we search for corresponding training answers on the training graph. For the validation queries, we use those queries that have different answers on the validation graph from answers on the training graph to evaluate the generalization ability. For the testing queries, we conduct a graph search on the testing graph with private edges for testing answers and we split the answers into private set and public set according to the definition of privacy risk query answers discussed in Section 4. We conduct a statistical analysis of the number of privacy risk answers for different types of complex queries of three knowledge graphs and the statistics are shown in Table 2.
@@ 6.3.1. Baselines
Our proposed P-NGDBs can be applied to various complex query encoding methods to provide privacy protections. We select three commonly used complex query encoding methods and compare the performance with and without P-NGDB’s protection. To better encode the numerical attribute values, we also use NRN (Bai et al., 2023a) in the query encoding process to handle number distributions. The detailed equations of the implementations of these backbones are given in Appendix A.
@@ 6.3.2. Evaluation Metrics
The evaluation consists of two distinct parts: reasoning performance evaluation and privacy protection evaluation. Following the previous work (Bai et al., 2023a), given a testing query $q$, the training, validation, and public testing answers are denoted as $M_{train}$, $M_{val}$, and $M_{test}$, respectively. We evaluate the quality of retrieved answers using Hit ratio (HR) and Mean reciprocal rank (MRR). HR@K metric evaluates the accuracy of retrieval by measuring the percentage of correct hits among the top K retrieved items. The MRR metric evaluates the performance of a ranking model by computing the average reciprocal rank of the first relevant item in a ranked list of results. we evaluate the generalization capability of models by calculating the rankings of answers that cannot be directly retrieved from an observed knowledge graph, which is $M_{test}/M_{val}$. For reasoning performance evaluation, higher metric values denote better retrieval quality. For privacy protection evaluation, we compute the metric of privacy-threatening answers as these answers cannot be inferred from the observed graphs. Because we want to obfuscate those answers, lower values denote stronger protection. We train all the models by using the training queries and private attributes and tune hyper-parameters using the validation queries. The evaluation is then finally conducted on the testing queries, including the evaluation of public answers for performance assessment and the evaluation of private answers for privacy assessment. The experiment results are reported on the testing public and private queries respectively.
@@
Table 4. The mean reciprocal ranking (MRR) variation for P-NGDBs on different types of queries. The performance ratio to the corresponding unprotected models is reported in parentheses.
@@ 6.5. Sensitivity Study
The privacy protection needs can be different in various knowledge graphs. We evaluate the impact of the privacy penalty coefficient $\beta$, A larger $\beta$ can provide stronger privacy protection. We select $\beta$ from $\{0.01,0.05,0.1,0.5,1\}$, and evaluate the retrieval accuracy of public answers and private answers, respectively. We evaluate GQE’s retrieval performance on the FB15K-N dataset and depict the MRR’s change under each $\beta$ compared to the unprotected GQE model. The results are shown in Figure 5. We know that the penalty can effectively control the level of privacy protection. For example, when $\beta=0.01$, the GQE model can achieve 97.4% MRR on public answers while only having 68.1% MRR on private answers. The flexible adjustment of the privacy coefficient can make P-NGDBs adapt to various scenarios with different privacy requirements. However, to meet higher privacy needs, P-NGDBs have to undertake much utility loss.
@@ 7. Conclusion
ACKNOWLEDGMENTS
The authors of this paper were supported by the NSFC Fund (U20B2053) from the NSFC of China, the RIF (R6020-19 and R6021-20) and the GRF (16211520 and 16205322) from RGC of Hong Kong. We also thank the support from UGC Research Matching Grants (RMGS20EG01-D, RMGS20CR11, RMGS20CR12, RMGS20EG19, RMGS20EG21, RMGS23CR05, RMGS23EG08).
Appendix A Instantiations of Query Encoders
In this section, we present detailed formulas for parameterizing various logical operations utilized in this paper. Additionally, we employ the following three query encoders: GQE, Q2B, and Q2P, as the foundational components of our approaches. Furthermore, to effectively handle numerical attribute values, we integrate the Number Reasoning Network (NRN) (Bai et al., 2023a) with these query encoding methods.
A.1. Graph Query Encoding (GQE)
In the GQE method (Hamilton et al., 2018), every query is represented as a singular vector. We denote the query embedding of each sub-query as $q_{i}\in\mathbf{R}^{k}$. In logical operations that involve multiple inputs, if there are multiple sub-queries, we denote them as $q_{i}^{1},q_{i}^{2},...,q_{i}^{n}$. Consistent with the paper, we employ the TransE variant of GQE for the instantiation of the relational projection in GQE. The specifics are outlined below. The equation for updating the query embedding in the GQE method is given by:
(10) $q_{i+1}=f_{P}(q_{i},r)=q_{i}+e_{r},$
where $e_{r}$ represents the relation embedding of $r$.
Additionally, for modeling the intersection operation, we utilize feed-forward functions and average pooling to implement a permutation-invariant neural network.
(11) $q_{i+1}=f_{I}(q_{i}^{1},...,q_{i}^{n})=W_{I}(\phi(\texttt{FFN}(q_{i}^{k}),\forall k\in\{1,2,...,n\})),$
where FFN is a feed-forward function, $\phi$ is an average pooling, and $W_{I}$ is parameter matrix. Following the setting in Q2B paper (Ren et al., 2020), we express the queries in disjunction normal form so that we can avoid the parameterization of union operations.
A.2. Box Embeddings (Q2B)
For the Q2B model, we implement the operations following the original paper. The box embedding can be separated into two parts, denoting the center and offsets, namely $q_{i}=[c_{i},o_{i}]$. Meanwhile, for each relation $r$, there is a corresponding box relation embedding $e_{r}=[c_{r},o_{r}]$. The relation projection $q_{i+1}=f_{P}(q_{i},r)$ is computed as follows
(12) $q_{i+1}=q_{i}+e_{r}=[c_{i}+c_{r},o_{i}+o_{r}]$
The intersection operations $q_{i+1}=f_{I}(q_{i}^{1},q_{i}^{1},...,q_{i}^{n})$ is computed as follows
(13) $c_{i+1}$ $=\sum_{k}a_{k}\cdot c_{i}^{k},\thickspace a_{k}=\frac{\texttt{exp}(\texttt{MLP}(q_{k}))}{\sum_{j}\texttt{exp}(\texttt{MLP}(q_{j}))}$
(14) $o_{i+1}$ $=\texttt{Min}(\{o_{i}^{1},o_{i}^{2},...,o_{i}^{n}\})\cdot\sigma(\texttt{DeepSets}(\{q_{i}^{1},q_{i}^{2},...,q_{i}^{n}\})),$
then the query embedding of $i+1$ step is $q_{i+1}=[c_{i+1},p_{i+1}]$. In this equation, $\cdot$ is the dimension-wise product, MLP is a multi-layer perceptron, and $\sigma$ is the sigmoid function. $\texttt{DeepSets}(\cdot)$ is the permutation-invariant deep architecture, which treats all the input sub-queries equally. The $\texttt{DeepSets}(\{x_{1},x_{2},...,x_{N}\})$ is implemented as $\texttt{MLP}(\frac{1}{N}\sum_{i}(\texttt{MLP}(x_{i})))$. Following the original paper (Ren et al., 2020), we use the disjunction normal form to avoid the direct parameterization of the union operation.
A.3. Particle Embeddings (Q2P)
In the Q2P embedding method (Bai et al., 2022), query embeddings are represented as a set of particle embeddings. Specifically, the query embedding can be represented as $q_{i}=[p_{i}^{1},p_{i}^{2},...,p_{i}^{k}]$, where $k$ denotes the number of particles in the embedding.
The relation projection $f_{P}$ is defined as follows: $q_{i+1}=f_{P}(q_{i},e_{l})$, where $q_{i}$ and $q_{i+1}$ represent the input and output particle embeddings, respectively.
Unlike the approach in (Bordes et al., 2013), where the same relation embedding $e_{l}$ is directly added to all particles in $q_{i}$ to model the relation projection, Q2P introduces multiple neuralized gates. These gates enable individual adjustment of the relation transition for each particle in $q_{i}$. The formulation is as follows:
(15) $Z=\sigma(W^{P}_{z}e_{l}+U_{z}q_{i}+b_{z}),$
(16) $R=\sigma(W^{P}_{r}e_{l}+U_{r}q_{i}+b_{r}),$
(17) $T=\phi(W^{P}_{h}e_{l}+U_{h}(R\odot q_{i})+b_{h}),$
(18) $A_{i}=(1-Z)\odot q_{i}+Z\odot T.$
Here, $\sigma$ and $\phi$ are the sigmoid and hyperbolic tangent functions, and $\odot$ is the Hadamard product. To allow information exchange among different particles, a scaled dot-product self-attention (vaswani2017attention) module is also incorporated,
(19) $q_{i+1}=\texttt{Attn}(W^{P}_{q}A^{T}_{i},W^{P}_{k}A^{T}_{i},W^{P}_{v}A^{T}_{i})^{T}.$
The parameters $W^{P}_{q},W^{P}_{k},W^{P}_{v}\in\mathbb{R}^{d\times d}$ are utilized to model the input Query, Key, and Value for the self-attention module called Attn. The intersection operation $f_{I}$ is defined on multiple sets of particle embeddings $\{q^{(n)}_{i}\}_{n=1}^{N}$. It outputs a single set of particle embeddings $q_{i+1}=f_{I}(\{q^{(n)}_{i}\}_{n=1}^{N})$. The particles from the $\{q^{(n)}_{i}\}_{n=1}^{N}$ are first merged into a new matrix $q_{i}=[q^{(1)}_{i},q^{(2)}_{i},...,q^{(N)}_{i}]\in R^{d\times NK}$, and this matrix $q_{i}$ serves as the input of the intersection operation. The operation updates the position of each particle based on the positions of other input particles. This process is modeled using scaled dot-product self-attention, followed by a multi-layer perceptron (MLP) layer. The equations for this process are as follows:
(20) $A_{i}=\texttt{Attn}(W^{I}_{q}q^{T}_{i},W^{I}_{k}q^{T}_{i},W^{I}_{v}q^{T}_{i})^{T},$
(21) $q_{i+1}=\texttt{MLP}(A_{i}).$
Here, $W^{I}_{q},W^{I}_{k},W^{I}_{v}\in\mathbb{R}^{d\times d}$ are the parameters for the self-attention layer. The MLP represents a multi-layer perceptron layer with ReLU activation. It is important to note that the parameters in the MLP layers of different operations are not shared.
A.4. Number Reasoning Network (NRN)
To better deal with numerical values in the queries, we incorporated NRN into these query encoding processes. Here are the implementation details.
A.4.1. Relational Projection and Logic Operations on Entities
In the number reasoning network, established query encoding methods are adopted as backbones for conducting relational projections and logic operations on entities. Three different encoding structures are utilized to encode the entity phase: query2box (Ren et al., 2020), GQE (Hamilton et al., 2018), and query2particles (Bai et al., 2022). Despite their differences, the query embeddings can be uniformly flattened to vectors, resulting in query encodings represented as $d$-dimensional vectors: $q_{i}\in\mathbb{R}^{d}$.
A.4.2. Attribute Projection
In the attribute projection step, given the query embedding $q_{i}$ of the entity variable at the $i$-th step and the attribute type $a$, we employ a neural structure $F_{p}$ to parameterize the distribution parameters for the $(i+1)$-th step:
(22) $\theta_{i+1}=F_{p}(q_{i},a),$
where $F_{p}$ is a neural structure whose weights are optimized during the learning process. A common parameterization for $F_{p}$ is the gated transition, inspired by the recurrent neural network unit. The parameterization equations are as follows:
(23) $\begin{split}&p_{i}=W^{p}_{p}q_{i}+b_{p}^{p}\\ &z_{i}=\sigma(W^{p}_{z}e_{a}+U^{p}_{z}p_{i}+b_{z}^{p}),\\ &r_{i}=\sigma(W^{p}_{r}e_{a}+U^{p}_{r}p_{i}+b^{p}_{r}),\\ &t_{i}=\phi(W^{p}_{h}e_{a}+U_{h}(r_{i}\odot p_{i})+b^{p}_{h}),\\ &\theta_{i+1}=(1-z_{i})\odot p_{i}+z_{i}\odot t_{i}.\end{split}$
where $e_{a}\in\mathbb{R}^{k}$ is the embedding used to encode the attribute type $a$. Here, $\sigma$ and $\phi$ represent the sigmoid and hyperbolic tangent functions, respectively, and $\odot$ denotes the Hadamard product. The parameter matrices are $W^{p}_{p}\in\mathbb{R}^{d\times k}$ and $W^{p}_{z},W^{p}_{r},W^{p}_{h},U^{p}_{z},U^{p}_{r},U^{p}_{h}\in\mathbb{R}^{k\times k}$.
A.4.3. Reverse Attribute Projection
To find entities that possess a numerical attribute value from a given set of values, the reverse attribute projection operation is employed to compute the query embedding for the resulting entity set:
(24) $q_{i+1}=F_{r}(\theta_{i},a),$
where $F_{r}$ is a neural structure with weights that are optimized during the learning process. The gated transition function in Equation (23) can also be used for reverse attribute projection. However, for this operation, the weights are substituted with a different set of parameters: $W^{r}_{p}$, $W^{r}_{z}$, $W^{r}_{r}$, $W^{r}_{h}\in\mathbb{R}^{k\times d}$, and $U^{r}_{z}$, $U^{r}_{r}$, $U^{r}_{h}\in\mathbb{R}^{d\times d}$.
A.4.4. Numerical Projection
The numerical projections are directly applied to the distribution parameters. Since the numerical constraint of $f$ can be an arbitrary binary function from $\mathbb{R}^{2}$ to ${0,1}$, we represent the projection operation using the function $F_{f}$ as follows:
(25) $\theta_{i+1}=F_{f}(\theta_{i}).$
A.4.5. Intersection and Union on Attribute Values
Suppose we have $n$ attribute variables $\theta_{i}^{1},\theta_{i}^{2},...,\theta_{i}^{n}$ as inputs for intersection and union operations. Their intersection and union can be expressed as follows:
(26) $\begin{split}\theta_{i+1}=F_{I}(\theta_{i}^{1},\theta_{i}^{2},...,\theta_{i}^{n}),\\ \theta_{i+1}=F_{U}(\theta_{i}^{1},\theta_{i}^{2},...,\theta_{i}^{n}),\end{split}$
where $F_{I}$ and $F_{U}$ are the functions that compute the intersection and union, respectively.
For the functions $F_{I}$ and $F_{U}$, we utilize the DeepSet (zaheer2017deep) functions to parameterize them. Specifically, we employ a self-attention network followed by a feed-forward network to implement the permutation-invariant DeepSet function. Given the input $\theta_{i}=[\theta_{i}^{1},\theta_{i}^{2},...,\theta_{i}^{n}]$, their formulas are as follows:
(27) $a_{i+1}$ $=\texttt{Attn}(W_{q}\theta_{i}^{T},W_{k}\theta_{i}^{T},W_{v}\theta_{i}^{T})^{T},$
(28) $\theta_{i+1}$ $=\texttt{MLP}(a_{i+1}),$
where $W_{q},W_{k},W_{v}\in\mathbb{R}^{k\times k}$ are parameters used to model the input Query, Key, and Value for the attention module Attn.